6. At the ends of the homogeneous rod supported on the point O, the forces F1 = 100N and F2 = 350N operate. Find the weight of the rod, knowing that it is in balance.


1
Expert's answer
2021-03-14T19:16:03-0400

For the equilibrium:


"W=F_1+F_2\\W=100+350=450\ N"